40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2536     Entry time: Wed Mar 10 17:20:26 2021
Author: Aidan 
Type: Update 
Category: 2um Photodiodes 
Subject: Pumped chamber down briefly - connected vacuum gauge to the computer 
  • I updated the SR560 settings for the AC Photodiode readout going into the ADC. They're now: AC coupled input, 2000x gain, low noise, 10kHz pole
  • Updated the "undoSR560" filterbank for the FM31 input channel to undo the SR560 settings (zpk([0.15], [0], 7.5E-5)
  • Unhooked oscilloscope from AC channel and switched SR560 to battery mode - no change in recorded spectra. I think I need to run the FEMTO preamp from a battery pack
  • Hooked up the pumping station to the chamber and went through the pumping cycle described below. No major issues.
  • Turned on the MKS vacuum gauge and hooked it into the serial port of QIL-WS1 to readback the vacuum pressure onto the DAQ. I want to transfer this across to QIl-NFS but there are no serial ports on that machine - will have to get a USB-Serial converter.
Attachment 1: IMG_1204.jpg  2.098 MB  | Hide | Hide all
Attachment 2: IMG_1200.jpg  2.540 MB  | Hide | Hide all
Attachment 3: IMG_1191.jpg  2.340 MB  | Hide | Hide all
Attachment 4: IMG_1190.jpg  2.823 MB  | Hide | Hide all
Attachment 5: IMG_1199.jpg  1.914 MB  | Hide | Hide all
ELOG V3.1.3-